OpenStudy (anonymous):

Verify: sin^4x+2sin^2xcos^2x+cos^4x=1

OpenStudy (anonymous):

Can you figure out some formula here???

OpenStudy (anonymous):

\[\sin^4(x) + \cos^4(x) + 2 \sin^2(x) \cdot \cos^2(x)\] \[\implies \color{red}{(\sin^2(x))^2 + (\cos^2(x))^2 + 2 \sin^2(x) \cdot \cos^2(x)}\]

OpenStudy (anonymous):

Now if you let : \[\sin^2(x) = a \quad and \; \quad \cos^2(x) = b\] then: \[\implies a^2 + b^2 + 2ab = (.....)^2 \quad \quad guess ...\]